OpenStudy (anonymous):

How do you set this up? The sum of two numbers is 25 and their difference is 7. Find the numbers.

OpenStudy (anonymous):

16 and 9

OpenStudy (anonymous):

I just guessed and checked.

OpenStudy (accessdenied):

a,b are two numbers. a + b = 25 (sum of two numbers, a and b) a - b = 7 (difference of two numbers)

OpenStudy (anonymous):

@AccessDenied nice!

OpenStudy (anonymous):

Oh ok then